if one student from group.B comes to group.A their strength becomes equal but if one student from group.A comes to group.B the strength of group B becomes double . How many students are present in both groups?
1
Expert's answer
2020-12-18T10:12:27-0500

Solution:

Given that "There are two student groups. They are Group A and Group B."

Let the number of students in group A be x and number of students in group B be y.


There are two conditions given:

(I) If one student from group B comes to group A their strength becomes equal.

⇒ x +1 = y - 1

⇒ x - y = -2 .......................................(i)


(II) If one students from group A comes to group B, the strength of group B becomes double.

⇒ 2(x -1) = y + 1

⇒ 2x - 2 = y + 1

⇒ 2x - y = 3........................................(ii)

Now we substract equation (ii) from (i), we get

⇒ x - y-(2x - y) = - 2-(3)

⇒ x - y - 2x + y = -2-3

⇒ - x = - 5

⇒ x = 5

Substitute x = 5 into equation (i), we get

⇒ x - y = - 2

⇒ 5 - y = - 2

⇒ - y = - 2 - 5

⇒ - y = - 7

⇒ y = 7


Therefore,

Number of students present in group A = 5 and

Number of students present in group B = 7